# Iterative Repair with Weak Verifiers for Few-shot Transfer in KBQA with Unanswerability

### v2

@@ Abstract
Real-world applications of KBQA require models to handle unanswerable questions with a limited volume of in-domain labeled training data. We propose the novel task of few-shot transfer for KBQA with unanswerable questions and contribute two new datasets for performance evaluation. We present FUn-FuSIC – a novel solution for our task that extends FuSIC-KBQA, the state-of-the-art few-shot transfer model for answerable-only KBQA. We first note that FuSIC-KBQA’s iterative repair makes a strong assumption that all questions are answerable. As a remedy, we propose Feedback for Unanswerability (FUn), which uses iterative repair using feedback from a suite of strong and weak verifiers, and an adaptation of self-consistency for unanswerabilty to better assess the answerability of a question. Our experiments show that FUn-FuSIC significantly outperforms suitable adaptations of multiple LLM-based and supervised SoTA models on our task, while establishing a new SoTA for answerable few-shot transfer as well.
@@
Riya Sawhney, Samrat Yadav, Indrajit Bhattacharya, Mausam Indian Institute of Technology, Delhi, KnowDis AI riya.sawhney@outlook.com, samratya23@gmail.com, indrajitb@gmail.com, mausam@cse.iitd.ac.in
@@ 1 Introduction
The semantic parsing formulation of the Knowledge Base Question Answering (KBQA) task takes as input a Knowledge Base (KB) and a natural language question, and outputs a logical form (or program) that produces the answer upon execution over the KB. KBQA has important real-world applications, which can require systems to be low resource (i.e., trained only with a few task-specific labeled examples), and robust, specifically the ability to identify questions that cannot be answered based on existing KB.
Traditional supervised models (e.g., Ye et al. (2022); Shu et al. (2022); Gu et al. (2023)) and even recent LLM few-shot in-context learning (FS-ICL) architectures Li et al. (2023); Nie et al. (2024) fall short in both aspects. Limited recent work has addressed these independently – in-domain methods for KBQA with unanswerability trained with large labeled data Patidar et al. (2023); Faldu et al. (2024), and FuSIC-KBQA for few-shot transfer assuming questions are answerable Patidar et al. (2024). No existing work simultaneously addresses both desiderata in a single KBQA system.
In response, we propose the novel task of few-shot transfer learning for KBQA with unanswerability. Specifically, the target domain may offer only a few labeled examples of answerable and unanswerable questions, while the source domain has thousands of labeled examples, but containing only answerable questions.
For KBQA transfer, FuSIC-KBQA uses a retrieve-then-generate framework: retrieval of relevant schema and KB snippets followed by an LLM-based generation and a subsequent iterative error guided repair. Specifically, multiple iterations of execution-guided feedback are run until a non-empty answer is obtained – this assumes that all questions are answerable. For our task, we first design FuSIC-KBQA-U, which, to address Unanswerability, eschews the now-inappropriate repair step, and modifies the LLM prompt to accommodate unanswerable questions, along with relevant in-context exemplars. Unlike in studies for unanswerability in general QA Slobodkin et al. (2023), this vanilla adaptation mostly generates incorrect logical forms for unanswerable questions.
@@
FUn-FuSIC’s verifiers consider both the logical form and the answer. For answers, non-emptiness check is now a weak verifier, given potentially unanswerable questions. For logical forms, we use strong verifiers to identify obvious syntactic and semantic errors. We also propose a novel verifier: non-equivalence of the original question and the back-translation of the logical form. This is weak, due to potential errors in back-translation as well as in equivalence classification.
Using such iterative weak verification based repair, FUn-FuSIC constructs a set of candidate logical forms. Simply selecting from this set the logical form with majority answer via self-consistency Wang et al. (2023) breaks down in the face of unanswerability. We introduce self-consistency for unaswerability, which additionally checks the likelihood of the majority answer, empty or otherwise, to produce the final output.
For evaluation in our novel task, we create two KBQA transfer datasets with unanswerable questions in the target. Our experiments show that FUn-FuSIC comprehensively outperforms different types of SoTA models suitably adapted for this task, including LLM-based and more traditional models. We further find that iterative repair of logical forms using weak verifiers holds promise for KBQA in general. Using experiments over benchmark datasets for answerable-only few-shot KBQA transfer, we show that the restriction of FUn-FuSIC for the answerable setting improves upon the known SoTA on the task.
In summary, our specific contributions are as follows. (a) We propose the problem of few-shot transfer for KBQA with unanswerability. (b) We present FUn-FuSIC that uses iterative repair with error feedback from a diverse suite of weak verifiers. (c) We create new datasets for the proposed task, which we make public. [^1] https://anonymous.4open.science/r/FUn-FuSIC-1704 (d) We show that FUn-FuSIC outperforms adaptation of SoTA KBQA models for our setting. (e) We also show that for standard answerable-only KBQA, FUn-FuSIC outperforms the corresponding SoTA model.
@@ 2 Related Work
For in-domain KBQA, unanswerability has recently been studied Patidar et al. (2023); Faldu et al. (2024).  Patidar et al. (2023) create the GrailQAbility dataset with different categories of unanswerability, and show the inadequacy of superficial adaptations of answerable-only KBQA models. RetinaQA Faldu et al. (2024) is the SoTA model for KBQA unaswerability. However, this also requires large volumes of training data.
For KBQA transfer Cao et al. (2022); Ravishankar et al. (2022), low-resource was originally not a focus. More recently, few-shot transfer for KBQA has been addressed Patidar et al. (2024). Its FuSIC-KBQA model uses a retrieve-then-generate framework with an LLM-based generation stage with iterative error feedback based repair.
However, this formulation assumes answerability of all questions, as described in Section 1.
@@
The extensive recent work on program self-repair using LLMs Olausson et al. (2024) also diverges strongly from our few-shot goal. It assumes availability of unit tests for every test instance. More importantly, as in the answerable KBQA literature, it assumes that a valid program exists for each specification.
To the best of our knowledge, ours is the first work to apply iterative weak repair for KBQA, and to use iterative repair for unanswerability.
@@ 3 Background & Problem Definition
Figure 1: Feedback with Unanswerability (FUn) and self consistency for Unanswerability (scUn) for a question $Q$ when executed over different KBs with $\leq 4$ iterations. $Q$ is answerable for KB3, but unanswerable for KB1 (schema incompleteness) and KB2 (data incompleteness). FUn iterations are shown using gray blocks, with Strong Verifiers in red and Weak Verifiers in green. (V1: Syntax Error omitted for brevity) scUn is shown using the blue block.
@@
A Knowledge Base (KB) $G$ consists of a schema and data. The schema consists of entity types (or classes) $T$ and binary relations $R$ defined over pairs of types. The data consists of entities $E$ as instances of types $T$, and triples or facts $F\subseteq E\times R\times E$. Given a target KB $G^{t}$ and a natural language question $q^{t}$, the goal is to generate a structured query or logical form $l^{t}$ (in a KB query language, such as SPARQL), which when executed over $G^{t}$ returns an answer $A^{t}$ (which is a set in general). Target few-shots $D^{t}$ containing tens of labeled training examples of questions and logical forms. A related source domain has a source KB $G^{s}$ (with its own types, relations, entities and facts), and a larger source training set $D^{s}$ with thousands of labeled training examples.
@@
Following Patidar et al. (2023), a question $q$ is answerable for a KB $G$ if it has a corresponding logical form $l$ which when executed over $G$ returns a non-empty answer $A$. A question is unanswerable, if it either (a) does not have a valid logical form for $G$ (schema-level unanswerability), or (b) it has a valid logical form $l$ for $G$, but $l$ returns an empty answer upon execution on $G$ (data-level unanswerability). Schema-level unanswerability arises due to missing types and relations, while missing entities and facts lead to data-level unanswerability. More details are in Appendix  A.1.3.
@@
In KBQA with unanswerability, given a question $q$, the model needs to output (a) a logical form $l$ and a non-empty answer A for answerable $q$, (b) $l=$ NK (No Knowledge) for schema-level unanswerable $q$, or (c) a valid logical form $l$ and $a=$ NA (No Answer) for data-level unanswerable $q$.
@@
We now define few-shot transfer learning for KBQA with unanswerability. A target question $q^{t}$ may be answerable or unanswerable due to missing schema or data in the target KB $G^{t}$. Target few shot examples $D^{t}$ contain both answerable and unanswerable questions of different categories. However, most KBQA datasets contain only answerable questions, we model the source training data $D^{s}$ as containing only answerable questions. Now the unanswerability gap between the source and the target also needs to be bridged. More details are in Appendix A.1.
@@ 4 Proposed Approach: FUn-FuSIC
Our proposed architecture FUn-FuSIC adapts FuSIC-KBQA for unaswerability. The high-level algorithm is described in Algo. 1. Preserving the retrieve-then-generate framework, the retrieval stage (line 2) performs KB retrieval for $q^{t}$ using a set $R$ of one or more supervised retrievers. Each retriever $R_{i}$ is source-trained and then target fine-tuned. The retrieval output $r$ of each $R_{i}$ consists of relevant schema elements (types and relations) for $q^{t}$, and data paths emanating from mentioned entities in $q^{t}$. The union of these, along with $q^{t}$, is fed to the generation stage, which uses prompting with an LLM $\mathcal{L}$ to generate logical forms using the target few shots $D^{t}$. More details of the retrieval stage are in the Appendix (Sec. A.6).
Algorithm 1 FUn-FuSIC$(q,G^{t},D^{t},R,V^{s},V^{w},\mathcal{L})$
1:$r=\{\}$
@@
3:$l=\mbox{PUn}(\mathcal{L},I,q,r,D^{t})$
4:$(e,l,A,L)=\mbox{FUn}(\mathcal{L},q,l,n,V^{s},V^{w},G^{t})$
5:if ($\mbox{{\bf if} }(e)\mbox{ {\bf return}}(l^{*},A^{*})$
6:else $\mbox{{\bf else} {\bf return}}(\mbox{scUn}(q,L,t,\mathcal{L}))$
FUn-FuSIC makes three modifications to the FuSIC-KBQA architecture. First, the LLM generation instruction is modified to admit the possibility of unanswerability, and the few shots are modified to include examples of unanswerable questions. However, this simple approach is error-prone. So, we bias the instruction towards one type of error. Specifically, when uncertain about answerability of the question, Prompting for Unanswerability (PUn) (line 3) instructs $\mathcal{L}$ to generate a (possibly incorrect) logical form instead of $l=\texttt{NK}$. The detailed prompt $I$ is in the Appendix (Sec. A.8).
We now come to the more significant modifications. First, $l^{(0)}$ is iteratively repaired using feedback as before, but this step is adapted for unaswerability. This iterative repair, which we name Feedback for Unaswerability (FUn) (line 4), either confidently outputs a single logical form $l$ (with corresponding answer $A$) (line 5) or generates a set $L$ of candidate logical forms, which is further analyzed. For this, we introduce self-consistency for Unaswerability (scUn). scUn checks the likelihood of the majority answer in $L$, empty or otherwise, to produce the final output. In the rest of this section, we describe FUn and then scUn. Fig. 1 illustrates flow of FUn and scUn using examples. A real example of FUn execution is in Sec. A.13.
@@
11:         if ($\mbox{{\bf if} }(e)\mbox{ }L=L\bigcup\{l\}$; $k++$
@@
13:     $l=Gen(\mathcal{L},I,q,F)$
14:end while
15:if ($\mbox{{\bf if} }(k=k_{2})\mbox{ {\bf return}}(\mbox{True},l,\mbox{Exec}(l,G),L)$
16:else $\mbox{{\bf else} {\bf return}}(\mbox{False},l,\{\},L)$
The FUn algorithm is described in Algo. 2 Starting with the initial logical form $l^{(0)}$, FUn performs at most $n$ verify-and-repair iterations to create a candidate set $L$ of probable logical forms. The example shows 3 FUn iterations for KB1 and KB2, and 2 iterations for KB3. In the $i^{th}$ iteration, FUn generates a new logical form $l^{(i)}$ by prompting $\mathcal{L}$ using $q^{t}$ and feedback $F$ received from checks in all previous iterations (line 13). $l^{(i)}$ goes through a sequence of verifications. FUn uses two sets of verifiers. The strong verifiers $V^{s}$ are guaranteed to be correct, while the weak verifiers $V^{w}$ are potentially erroneous. The specific verifiers that we use in this paper are defined later in the section. A template-based feedback string $f$ is appended to the generation prompt for $l^{(i+1)}$ based on the specific verifier that $l^{(i)}$ failed. If $l^{(i)}$ fails a strong verifier, it is rejected (line 6). In the example, this happens for all three KBs in iteration 1. If $l^{(i)}$ passes all checks, strong and weak (line 15), FUn terminates by outputting $(l=l^{(i)},A=A^{(i)})$, where $A^{(i)}$ is the answer obtained by executing $l^{(i)}$. This happens in iteration 3 for KB3. Otherwise, if $l^{(i)}$ passes at least one weak verifier but not all, it is added to candidate logical form set $L$ (line 11). This happens for iterations 2, 3 and 4 for KB1 and KB2.
@@
FUn uses a suite verifiers, categorized as strong ($V^{s}$) and weak ($V^{w}$). These may be syntactic, semantic, or execution-based, defined using simple rules or complex LLM functions over $l$, $q$ and $G$. Note that unlike unit tests in program synthesis, the verifiers do not have knowledge of the gold logical form, the gold answer or answerability of the question.
We now briefly describe the specific verifiers that we use for this paper. Additional details about the verifiers are in the Appendix (Sec. A.8.1 and Sec. A.2). Note however that FUn is a framework that is equally capable of working with a wholly different suite of verifiers.
@@
As in FuSIC-KBQA, this verifier executes the logical form $l$ over $G$ and checks for syntax error. This is a strong check: a valid logical form cannot have syntax error.
@@
This verifier checks if $l$ is semantically equivalent to the original natural language question $q$. In our example, LF3 for KB1 disagrees with Q. This is a weak verifier. First, $q$ may not have any equivalent logical form for $G$ due to intrinsic ambiguities. For example, $q$ mentions a person from a country, when $G$ has the relations born in and works in between these types. Secondly, this is implemented as a probabilistic classifier that can err. We define equivalence check between $l$ and $q$ using a novel multi-stage LLM pipeline, involving naturalization of $l$ to $l^{n}$, back-translation of $l^{n}$ to natural language question $q^{b}$ and semantic equivalence check between $q$ and $q^{b}$. More details are in Sec.A.2.
@@
Unless some logical form passes all checks and is therefore returned (Algo. 2 line 6), FUn constructs a candidate set $L$ of logical forms that are potentially flawed but not certainly so. For our specific suite of weak verifiers, $l^{(i)}$ is added to $L$ if it passes one of V4b ($A$ is non-empty) as for LF3 for KB1, or V3 ($l^{i}$ is equivalent to $q$), as for LF3 for KB2.
Self Consistency for Unanswerability:
If FUn fails to generate a single confident logical form at the end of $n$ iterations, as for KB1 and KB2, we use scUn to decide if the best candidate $l^{*}\in L$ has sufficient confidence. If scUn detects sufficient confidence, as for KB2, it outputs ($l=l^{*}$, $A=A^{*}$), where $A^{*}$ is the answer obtained upon executing $l^{*}$ and may be NA. If on the other hand, scUn decides insufficient confidence, as for KB1, it outputs ($l=$ NK, $A=$ NA). For identifying the consensus choice from $L$, one possibility is self-consistency Wang et al. (2023); Chen et al. (2023a) i.e., obtaining the answer for each $l\in L$, and returning those with the most common answer. Self-consistency requires some answer to accumulate enough probability by aggregation over reasoning paths. However, for unanswerable questions, no single answer accumulates sufficient probability, and self-consistency returns some low probability answer.
To address this, scUn first identifies via execution the most popular non-empty answer $A^{*}$ among logical forms in $L$, and decides using threshold $t$ if it has enough supporters in $L$ (we use $t=\lfloor\frac{|L|}{2}\rfloor$). If so, scUn uses LLM prompting to select the most appropriate supporting logical form $l^{*}\in L$ considering $q$, and outputs ($l=l^{*}$, $A=A^{*}$).
However, for KB1 in the example, the 3 logical forms among the candidates have 3 different answers, and therefore ($\lfloor\frac{|L|}{2}\rfloor=1$) no consensus emerges for non-empty answers. Then scUn considers logical forms from $L$ that agree on $A=$ NA. If there are multiple such candidates, scUn selects the most suitable candidate $l^{*}$, again using LLM prompting, and outputs ($l=l^{*}$, $A=$ NA). If there is no such candidate, scUn outputs ($l=$ NK, $A=$ NA). For KB2 in the example, scUn selects LF3 – the only logical form with empty answer. Further details on scUn are in the Appendix (Sec. A.14). .

### v3

@@ Abstract
Real-world applications of KBQA require models to detect different types of unanswerable questions with a limited volume of in-domain labeled training data. We propose the novel task of few-shot transfer for KBQA with unanswerable questions. The state-of-the-art KBQA few-shot transfer model (FuSIC-KBQA) uses an iterative repair strategy that assumes that all questions are answerable. As a remedy, we present FUn-FuSIC – a novel solution for our task that extends FuSIC-KBQA with Feedback for Unanswerability (FUn), which is an iterative repair strategy for answerable as well as unanswerable questions. FUn uses feedback from a suite of strong and weak verifiers, and an adaptation of self-consistency for unanswerability for assessing answerability of questions. Our experiments show that FUn-FuSIC significantly outperforms suitable adaptations of multiple LLM-based and supervised SoTA models on our task, while establishing a new SoTA performance for answerable few-shot transfer as well. We have made datasets and other resources publicly available [^1] https://github.com/dair-iitd/FUn-FuSIC
@@
Riya Sawhney‡,§, Samrat Yadav, Indrajit Bhattacharya, Mausam Indian Institute of Technology, Delhi, §Graviton Research Capital, KnowDis AI riya.sawhney@outlook.com, samratya23@gmail.com, indrajitb@gmail.com, mausam@cse.iitd.ac.in
@@ 1 Introduction
The semantic parsing formulation of the Knowledge Base Question Answering (KBQA) task takes as input a Knowledge Base (KB) and a natural language question, and outputs a logical form (or program) that produces the answer upon execution over the KB. KBQA has important real-world applications, which require KBQA systems to be low-resource (i.e., trained only with a few task-specific labeled examples), and robust, specifically able to identify questions that cannot be answered from the KB.
Traditional supervised models (e.g., Ye et al. (2022); Shu et al. (2022); Gu et al. (2023)) and even recent LLM few-shot in-context learning (FS-ICL) architectures Li et al. (2023); Nie et al. (2024) for KBQA fall short in both aspects. Limited recent work has addressed these independently – in-domain methods for KBQA with unanswerability trained with large labeled data Patidar et al. (2023); Faldu et al. (2024), and FuSIC-KBQA for few-shot transfer assuming answerable questions Patidar et al. (2024). No existing single KBQA model simultaneously addresses both desiderata.
In response, we propose the novel task of few-shot transfer learning for KBQA with unanswerability. Specifically, the target domain has only a few labeled examples of answerable and unanswerable questions, while the source domain has thousands of labeled examples, but containing only answerable questions.
For few-shot KBQA transfer, FuSIC-KBQA uses a retrieve-then-generate framework: retrieval of relevant schema and KB snippets followed by an LLM-based generation and a subsequent iterative execution-error-guided repair. Specifically, multiple feedback-guided repair iterations are executed, checking emptiness of answers obtained by executing the generated program as indication of correctness, until a non-empty answer is obtained. This naturally fails when questions are allowed to be unanswerable.
A simple-fix for addressing Unanswerability (FuSIC-KBQA-U) is to drop the inappropriate repair step, and modify the LLM prompt to accommodate unanswerable questions, along with relevant in-context exemplars. Unlike in studies for unanswerability in general QA Slobodkin et al. (2023), we found that FuSIC-KBQA-U mostly generates incorrect logical forms for unanswerable questions.
@@
FUn-FuSIC’s verifiers consider both the logical form and the answer. For answers, non-emptiness check is now a weak verifier, given potentially unanswerable questions. For logical forms, we use strong verifiers to identify obvious syntactic and semantic errors. We also propose a novel verifier involving a 3-component LLM-based pipeline: non-equivalence of the original question and the back-translation of the logical form. This verifier is also weak, due to potential errors in back-translation as well as in equivalence classification. Using such iterative strong and weak verification based repair, FUn-FuSIC constructs a set of candidate logical forms. For selecting the consensus logical form from this set, using the majority answer as in self-consistency Wang et al. (2023) breaks down in the face of unanswerability. We introduce self-consistency for unanswerability, which assesses the likelihood of the majority answer, empty or otherwise, to select the consensus logical form.
Since no datasets exist for our novel task, we create two new datasets for KBQA transfer with unanswerable questions. Our experiments show that FUn-FuSIC comprehensively outperforms different categories of SoTA models suitably adapted for this task, including LLM-based and more traditional models. We further find that iterative repair of logical forms using weak verifiers holds promise for even for KBQA with only answerable questions. Using experiments over benchmark datasets for this task, we show that the restriction of FUn-FuSIC for the answerable setting improves upon the SoTA model for the task.
In summary, our specific contributions are as follows. (a) We propose the problem of few-shot transfer for KBQA with unanswerability. (b) We present FUn-FuSIC that uses iterative repair with error feedback from a diverse suite of strong and weak verifiers. (c) We create new datasets for the proposed task, which we make public. (d) We show that FUn-FuSIC outperforms adaptations of SoTA KBQA models for this new task. (e) We also show that even for answerable-only KBQA, FUn-FuSIC outperforms the corresponding SoTA model.
@@ 2 Related Work
For in-domain KBQA, unanswerability has recently been studied Patidar et al. (2023); Faldu et al. (2024).  Patidar et al. (2023) create the GrailQAbility dataset with different categories of unanswerability, and show the inadequacy of superficial adaptations of answerable-only KBQA models. RetinaQA Faldu et al. (2024) is the SoTA model for KBQA unanswerability. However, this also requires large volumes of training data.
For KBQA transfer Cao et al. (2022); Ravishankar et al. (2022), low-resource was originally not a focus. More recently, few-shot transfer for KBQA has been addressed by FuSIC-KBQA Patidar et al. (2024). FuSIC-KBQA uses a retrieve-then-generate framework with an LLM-based generation stage with iterative error feedback based repair. However, this formulation assumes answerability of all questions.
@@
FUn’s iterative repair idea may be useful in other natural language to program generation tasks where non-existence of a program with the required specification has not been studied to the best of our knowledge, such as NL-to-SQL  Dong et al. (2023); Pourreza and Rafiei (2023) and program self-repair using LLMs Olausson et al. (2024); Madaan et al. (2023); Grattafiori et al. (2024).
@@ 3 Background & Problem Definition
Figure 1: Feedback with Unanswerability (FUn) and self consistency for Unanswerability (scUn) for a question when executed over different KBs with 4$\leq$ 4≤ 4 iterations. The question is answerable for KB3, but unanswerable for KB1 (schema incompleteness) and KB2 (data incompleteness). In the KB depictions, the top graph represents the schema with different node colors for different entity types, and the bottom graph represents the data. (Names of different (real) books related to the author entity in the question are abbreviated in the data graph.) FUn iterations are shown in the gray blocks, with Strong Verifiers named using red text and Weak Verifiers using green text and $i$ denotes the iteration number. The outcome of verification is denoted as True or False, and the non-empty answer is shown when V4b returns False. The Syntax Error Verifier (V1) is omitted for brevity. scUn is shown in the blue blocks. The candidate logical forms ($L$) for sCun are shown at the top. For Non-empty Ans and Empty Ans agreement checks for, the outcome is either ‘Consensus’ or ‘No consensus’. At the end, FUn-FuSIC ‘Return’s a logical form (possibly NK) and an answer (possibly NA).
@@
A Knowledge Base (KB) $G$ consists of a schema and data. The schema consists of entity types (or classes) $T$ and binary relations $R$ defined over pairs of types. The data consists of entities $E$ as instances of types $T$, and triples or facts $F\subseteq E\times R\times E$. Given a target KB $G^{t}$ and a natural language question $q^{t}$, the basic KBQA task is to generate a structured query or logical form $l^{t}$ (in a KB query language, such as SPARQL), which when executed over $G^{t}$ returns an answer $A^{t}$ ($A^{t}\subset E$ in general). Other than SPARQL  Patidar et al. (2024), niche languages such as s-expressions  Li et al. (2023); Gu et al. (2023) are commonly used for logical forms in KBQA. In supervised in-domain KBQA, the target has large volumes of labeled training examples of questions and associated logical forms. For few-shot in-domain KBQA in contrast, target few-shots $D^{t}$ contain tens of labeled training examples. In few-shot transfer learning for KBQA Patidar et al. (2024), a related source domain has a source KB $G^{s}$ (with its own types, relations, entities and facts), and a larger source training set $D^{s}$ with thousands of labeled training examples.
@@
Following Patidar et al. (2023), a question $q$ is answerable for a KB $G$ if it admits a corresponding logical form $l$ which when executed over $G$ returns the ideal non-empty answer $A$. A question is unanswerable if it either (a) does not have a valid logical form for $G$ (schema-level unanswerability), or (b) it has a valid logical form $l$ for $G$, but $l$ returns an empty answer upon execution on $G$, different from the ideal non-empty answer (data-level unanswerability). Schema-level unanswerability arises due to missing types and relations, while missing entities and facts lead to data-level unanswerability. However, absence in $G$ of any entity mentioned in the question is categorized as schema level unanswerability, since it invalidates the logical form. More details are in Appendix  A.1.2.
@@
In KBQA with unanswerability Patidar et al. (2023); Faldu et al. (2024), given a question $q$, the model needs to output (a) a logical form $l$ and a non-empty answer A for answerable $q$, (b) $l=$ NK (No Knowledge) for schema-level unanswerable $q$, or (c) a valid logical form $l$ and $a=$ NA (No Answer) for data-level unanswerable $q$. In the supervised in-domain setting, this task involves large volumes of labeled training questions, containing both answerable and unanswerable, for the target.
@@
We now define our problem of interest: few-shot transfer learning for KBQA with unanswerability. A target question $q^{t}$ may be answerable or unanswerable due to missing schema or data in the target KB $G^{t}$. Target few-shot examples $D^{t}$ contain both answerable and unanswerable questions of different categories. The source training data $D^{s}$ has large volumes of labeled training data. Considering real world constraints, where most KBQA datasets contain only answerable questions, we assume that $D^{s}$ contains only answerable questions. Compared to the earlier few-shot KBQA transfer task definition, now there is additionally an unanswerability mismatch between the source and the target distribution. More details are in the Appendix (Sec.  A.1).
@@ 4 Proposed Approach: FUn-FuSIC
Our proposed model FUn-FuSIC preserves the basic architecture of FuSIC-KBQA and adapts its iterative repair strategy for unanswerability. The high-level algorithm is described in Algo. 1. (Since the algorithms are not specific to the transfer task, we use $q$ instead of $q^{t}$ for brevity.) Preserving the retrieve-then-generate framework of FuSIC-KBQA, the retrieval stage (line 2) performs KB retrieval for $q^{t}$ using a set $R$ of one or more supervised retrievers. Each retriever $R_{i}$ is source-trained and further target fine-tuned if required. The retrieval output $r$ of each $R_{i}$ consists of relevant schema elements (types and relations) for $q^{t}$, and data paths emanating from mentioned entities in $q^{t}$. The union of these, along with $q^{t}$, is fed to the generation stage, which uses prompting with an LLM $\mathcal{L}$ to generate logical forms using the target few-shots $D^{t}$. More details of the retrieval stage are in the Appendix (Sec. A.8.1).
@@
[…]
6:else return $\mbox{scUn}(q,L,\mathcal{L})$
FUn-FuSIC differs from FuSIC-KBQA in its iterative repair strategy in lines 3, 4 and 6 of Algo. 1. The LLM generation instruction is modified to admit the possibility of unanswerability, and the few shots are modified to include examples of unanswerable questions. However, this simple approach is error-prone. So, we bias the instruction towards one type of error. Specifically, when uncertain about answerability of the question, Prompting for Unanswerability (PUn) (line 3) instructs $\mathcal{L}$ to generate a (possibly incorrect) logical form instead of $l=\texttt{NK}$. The detailed prompt $I$ is in the Appendix (Sec. A.11.1).
We now come to the more significant modifications. First, $l^{(0)}$ is iteratively repaired using feedback as before, but this step is adapted for unanswerability. This iterative repair, which we name Feedback for Unanswerability (FUn) (line 4), either confidently outputs a single logical form $l$ (with corresponding answer $A$) (line 5) or generates a set $L$ of candidate logical forms, which is further analyzed for a consensus logical form and answer. For this, we introduce self-consistency for Unanswerability (scUn) (line 6). scUn assesses the likelihood of the majority answer in $L$, empty or otherwise, to produce the final output. In the rest of this section, we describe FUn and then scUn. Fig. 1 illustrates flow of FUn and scUn using examples. A real example of FUn execution is in Sec. A.12.
@@
11:   $\mbox{{\bf if} }(e)\mbox{ }L=L\bigcup\{l^{(i)}\}$; $k++$
@@
13:  $i=i+1$
14:  $l^{(i)}=Gen(\mathcal{L},I,q,F)$
[…]
16:end while
[…]
The FUn algorithm is described in Algo. 2 Starting with the initial logical form $l^{(0)}$, FUn performs at most $n$ verify-and-repair iterations to create a candidate set $L$ of probable logical forms. Fig. 1 shows 3 FUn iterations for KB1 and KB2, and 2 for KB3. In the $i^{th}$ iteration, FUn generates a new logical form $l^{(i)}$ by prompting $\mathcal{L}$ using $q^{t}$ and feedback $F$ received from checks in all previous iterations (line 13). $l^{(i)}$ goes through a sequence of verifications. FUn uses two sets of verifiers. The strong verifiers $V^{s}$ are guaranteed to be correct, while the weak verifiers $V^{w}$ are potentially erroneous. $k_{1}$ and $k_{2}$ denote the total number of strong and weak verifiers respectively. The specific verifiers that we use in this paper are defined later in the section. A template-based feedback string $f$ is appended to the generation prompt for $l^{(i+1)}$ based on the specific verifier that $l^{(i)}$ failed. If $l^{(i)}$ fails a strong verifier, it is rejected (line 6). In the example, this happens for all three KBs in iteration 1. If $l^{(i)}$ passes all checks, strong and weak (line 15), FUn terminates by outputting $(l=l^{(i)},A=A^{(i)})$, where $A^{(i)}$ is the answer obtained by executing $l^{(i)}$ (denoted $\mbox{X}(l^{(i)},G)$). This happens in iteration 3 for KB3. Otherwise, if $l^{(i)}$ passes at least one weak verifier but not all, it is added to candidate logical form set $L$ (line 11). This happens for iterations 2, 3 and 4 for KB1 and KB2.
@@
FUn uses a suite of verifiers, categorized as strong ($V^{s}$) and weak ($V^{w}$). These may be syntactic, semantic, or execution-based, defined using simple rules or complex LLM functions over $l$, $q$ and $G$. Note that unlike unit tests in program synthesis, the verifiers do not have knowledge of the gold logical form, the gold answer or answerability of the question.
We now briefly describe the specific verifiers that we use for this paper. Additional details about the verifiers are in the Appendix (Sec. A.11.2 and Sec. A.3). Note that FUn is a framework that is capable of working with a wholly different suite of meaningful verifiers.
@@
As in FuSIC-KBQA, this verifier executes the logical form $l$ over $G$ and checks for syntax error. This is a strong check — a valid logical form cannot have syntax error.
@@
This verifier checks if $l$ is semantically equivalent to the original natural language question $q$. In Fig. 1, LF3 for KB1 disagrees with $q$. This is a weak verifier. First, $q$ may not have any equivalent logical form for $G$ due to intrinsic ambiguities even when it is answerable. For example, $q$ mentions a person from a country, when $G$ has the relations born in and works in between these types. Secondly, this verifier is a probabilistic classifier that naturally makes occasional mistakes. We define equivalence check between $l$ and $q$ using a multi-stage LLM pipeline, involving naturalization of $l$ to $l^{n}$, back-translation of $l^{n}$ to natural language question $q^{b}$ and semantic equivalence check between $q$ and $q^{b}$. More details are in the Appendix (Sec.A.3).
@@
Unless some logical form passes all checks and is therefore returned (Algo. 2 line 15), FUn constructs a candidate set $L$ of logical forms that are potentially flawed but not certainly so. For our specific suite of weak verifiers, $l^{(i)}$ is added to $L$ if it passes one of V4b ($A$ is non-empty) as for LF3 for KB1, or V3 ($l^{i}$ is equivalent to $q$), as for LF3 for KB2.
Self Consistency for Unanswerability (scUn):
Given a candidate set $L$ of logical forms and a question $q$, scUn assesses if the best candidate $l^{*}\in L$ has sufficient confidence. If so, it outputs ($l=l^{*}$, $A=A^{*}$), as for KB2, $A^{*}$ being the answer from executing $l^{*}$ (may be NA). Otherwise, scUn outputs ($l=$ NK, $A=$ NA), as for KB1. For identifying the consensus choice from $L$, one possibility is self-consistency (sc) Wang et al. (2023); Chen et al. (2023a) that considers the answer for each $l\in L$, and returns those with the most common answer. This requires some answer to accumulate enough probability by aggregation over reasoning paths. However, for unanswerable questions, no single answer accumulates sufficient probability, and sc returns some low probability answer.
To address this, scUn first identifies via execution the most popular non-empty answer $A^{*}$ among logical forms in $L$, and decides using a threshold $t$ if it has enough supporters in $L$ (we use $t=\lfloor\frac{|L|}{2}\rfloor$). If so, scUn uses LLM prompting to select the most appropriate supporting logical form $l^{*}\in L$ considering $q$, and outputs ($l=l^{*}$, $A=A^{*}$). However, for KB1, the 3 logical forms among the candidates have 3 different answers, and therefore no consensus emerges ($\lfloor\frac{|L|}{2}\rfloor=1$). Here, scUn considers logical forms from $L$ that agree on $A=$ NA. If there are multiple such candidates, scUn selects the most suitable candidate $l^{*}$, again using LLM prompting, and outputs ($l=l^{*}$, $A=$ NA). If there is no such candidate, scUn outputs ($l=$ NK, $A=$ NA). For KB2 in the example, scUn selects LF3 – the only logical form with empty answer. Further details on scUn are in the Appendix (Sec. A.2).